2) Clifftown’s Parks and Recreation Department is introducing a new summer day camp for children in elementary school, providing programming from 8:00 a.m. to 5:00 p.m. Monday through Friday. The proposed day camp program is 10 weeks long and is planned for 50 children. Fixed costs include staff, equipment, and facilities costs, and the total fixed cost is estimated at $15,000 for the 10 weeks. The camp will serve lunch and snacks each weekday at a weekly cost of $15 per child. How much should the Parks and Recreation Department charge per child for the program to break even? *** BREAK EVEN FORMULA**** What is the FC -FV / VR - VC
Answer #1

Breakeven is a situation which derives no profit and no loss means our sales should be equal to Variable cost and Fixed Cost

Variable Cost VC = Cost of snacks per weekday per child * no. of children * no. of days in a week * no. of weeks

= 15*50*5*10 = 37500

Fixed Cost FC = 15000

Total Cost = 15000 + 37500 = 52500

How much to charge for Breakeven per child = 52500/50 = $1050